OpenStudy (anonymous):

Simplify the expression: [1/ (3+x) - 1/3] / x

OpenStudy (anonymous):

-1/(3(x+3))

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

If so how did you get that?

OpenStudy (anonymous):

so 1/(3+x) - 1/3 = 3/(3x+9) - (3+x)/(3x+9) = (3-(3+x))/(3x+9) Thus (-x/(3x+9))/x = -1/(3x+9)=-1/(3(x+3))

OpenStudy (anonymous):

thanks for teh solution